OpenStudy (anonymous):

Let f(x) = sin-1x. Then, the domain of f(x) is [-1,1] and the range is all real numbers. True or False

OpenStudy (anonymous):

Is it false?

OpenStudy (mathstudent55):

The range of sin x is [-1, 1] That means the domain of \(\sin ^{-1} x \) is [-1, 1]. The domain is correct.

OpenStudy (anonymous):

so True?

OpenStudy (mathstudent55):

The domain is correct, but the range is not. The range of the inverse sine function is limited to \([ -\dfrac{\pi}{2}, \dfrac{\pi}{2} ]\) The range is incorrect, so the answer must be false.

OpenStudy (anonymous):

Thanks:)

OpenStudy (mathstudent55):

You're welcome.